# Robust Deep Network Learning of Nonlinear Regression Tasks by Parametric Leaky Exponential Linear Units (LELUs) and a Diffusion Metric

### v1

@@ 4.1.3 Exponential Function and a Single Neuron
The previous problem is reduced to a set of three training points, $x_{i}=(-1,0,1)$, which a single neuron shall learn. This exercise can be performed by hand and two of the three points can be exactly matched. The obtained neuron parameters are used to infer a function as shown in Fig. 7, plotted along the ac.f. derivatives and training errors.
@@
From the perspective of learning nonlinear regression tasks, such weak signal is less then ideal. The gradient and Hessian of the loss can be computed by hand for this simple problem. This operation shows that the weak ac.f.’s produce a vanishing loss gradient and higher conditioned Hessians, which is indicative of the tendency to poor learning and overfitting. The LELU on the other hand produces a clear gradient and a lower conditioned Hessian. With this simple exercise, the author intends to anecdotally show that the higher “flexibility” of the ac.f. is incurring in a weak signal for subsequent neurons and is a potential cause of instability showing as overfitting.
@@ 4.1.4 Other Findings with the Unidimensional problems
It is interesting to note that some ac.f.’s struggle around $x=0$ in both case studies, even when this point is not a training data. A thorough explanation is challenging for a large model [16], but the author believes this can be partially understood by a similar analysis of the gradient as in the previous subsection. The loss-gradient is proportional to the summation of the gradient of the ac.f. times the training point coordinate. If the training point coordinate tends to zero, or adjacent training points cross $x=0$, this can further induce the loss gradient summation to vanish or “flicker”, and to destabilize the learning on an already unstable model setup. This behaviour is confirmed with a simple shift in $x$ by two units, as shown in Fig. 8, in which the local instability around $x=0$ is removed for the unstable models. Note that SiLU actually lost performance and presents overfitting, opposed to the behaviour shown in Fig. 5a, which corroborates the borderline stability of the model.
Another observation is that discontinuous ac.f.’s generate discontinuous trained models in both exercises. In these and some other setups not shown in here, the discontinuous ac.f.’s rather often default to piecewise linear connection between the training points. Even when this was not the case, the trained model is still discontinuous, as seen in Figs. 5, 6 and 8.
The adopted canonical problems indicate that classical ac.f.’s lack robustness with the problem setup for nonlinear regressions tasks when trained by large models. The proposition of the objectives in Sect. 2 for the LELU construction aims at mitigating the reported issues.
4.1.5 Other Approaches to Decrease Overfitting
By assuming that the author has been successful in the beginning of this section to justify the most obvious criticism against these exercises, i.e., the adopted model sizes, other strategies also merit further discussions. Known strategies to deal with overfitting, namely, regularisation and smaller learning rates for longer epochs were also evaluated. The results are not shown for the sake of conciseness since the outcomes are as expected, i.e., a decrease in the overfitting is achieved however at the expense of considerably higher training MAEs and/or computational resources. Both L1 and L2 regularisations, with constant strength $41e-$ 4, are able to reduce or remove (L1 regularisation) overfitting, however at the expense of $10x$ or larger training MAEs. In terms of learning rates, an example with the ELU required reduction of half of the initial learning rate, and $10x$ longer training epochs.
The LELU is able to achieve adequate results without those penalties. It can robustly learn the nonlinear regression task without memorising the training data. This behaviour is further confirmed on the more complex model discussed in what follows.
@@ 4.2 Electric Motor Model
The dataset is normalised so as to keep unitary mesh spacing in all dimensions to avoid cell skewness (which also favours the computation of the diffusion differentials), and kept in the positive range to support the more sensitive ac.f.’s. For the current study the trained data must exactly reproduce the noise-free training information [1]. A noiseless training is generally more challenging and is an additional reason for this choice in the current exercise.
A numerical experiment with the implemented [^2] Source files available in http://github.com/ebiga/gpr_aniso_trials NN with combinations of depth $n=[6,7,8]$ and width $m=[80,120,240,360]$ (fixed per layer) is executed. The model is trained for 9000 epochs with batch size 35. The learning rate is scheduled in the range $[1e-3,1e-5]$ ($45e-$ 4 for the models with width $360$ and the $8\times 240$ case) with factor 0.5, patience 250, and cooldown 50.
@@
Figure 10: Trained model losses for the electric motor.
Figure 10 shows the MAE and diffusion metrics for the different model sizes (indexed by $m^{n}/1e12$ in the plots). The tendency to overfit shows as a lack of “convergence” of the MAE with the increase in the model size. For this complex problem the mean errors do not show large differences between the “good” and “bad,” so visual inspection is also necessary. Nonetheless, it can be seen that both metrics are lower with the LELU and are also more robust to model size. SiLU and Softplus do not present any favourable behaviour as partially seen with the canonical problems. The conclusions above are corroborated by the slice plots in Figs. 11 and 12.
@@
Figure 11 shows slices along $param2$, which is the fast-changing variable. The tanh and ReLU results are not shown due to the poor performance. The region $param2=[0,2]$, where the highest slopes are seen, presents a tough task for the ac.f.’s. Considerable “localised” overfitting is observed for the ELU, SiLU and Softplus. Interestingly, the less flexible leaky ac.f.’s show lower overfitting content and also less dependency on the model size. This nice behaviour prompted the current effort into assembling this property into a new ac.f. The Leaky ReLU results are “kinky” though and such behaviour is not ideal. The proposed LELU with $\beta=0.4$ shows a robust performance, indicating that the lower flexibility and smoothness are indeed advantageous.
@@
Figure 12 shows similar slices along the tight- and slow-changing $param3$. This slice allows for interesting behaviours to be more closely assessed, i.e., the spread of the ac.f. with different model sizes and the capacity to code small values. The less flexible leaky ac.f.’s clearly show tighter and more representative capacity. SiLU and Softplus are also able to demonstrate a tighter spread, albeit with an offset error to the training data, whereas ELU is generally poorly behaved.
By inspection of ac.f. graphs in Fig. 1, it can be seen that both SiLU and Softplus present a wider range of slowly changing gradient, or in other words, ELU has a sharper gradient change. In the SiLU case, the gradient is actually never zero in the range plotted in Fig. 1. This observation has further prompted the effort into coding this property into the LELU. As seen in Fig. 1, this ac.f. slowly changes slopes and is therefore smoother than other ac.f.’s, thereby enabling a stronger a priori fitness for generalisable nonlinear regressions.
@@
For this motor model the LELU with $\beta=0.4$ provides the most agreeable results. Upon careful visual inspection of each NN result (see Fig. 13), depth 6 is not enough to capture all features, whereas depth 7 is able to do so but the MAE still indicates some distance (error) to the training points. Generally at least width 120 is necessary for acceptable results.
@@
It is worth revisiting the discussion in Sect. 4.1.5. Except from reducing the model size, which is in fact hopefully clear now the need for such larger models, the other techniques to control overfitting also apply to the motor model. As shown in Fig. 14, both L1 and L2 regularisations (strength $41e-$ 4) reduce the overfitting behaviour of the SiLU, at the expense of representation of training points. The MAE increase by more than a factor of 6 in these cases and the performance does not improve over the LELU ($\beta=0.4$). Another strategy is to implement a change of variables, for instance a power correction $\tilde{y}=y^{0.1}$, to reduce the large differences in the max and min values. A model trained on this function learns faster and easier, however the resulting errors when brought back to the original scale are also very large, which defeats the purpose. As already observed with the canonical problems, the LELU seems to provide adequate results without those drawbacks and extra steps.
@@ 5 Concluding Remarks
A parametric activation function for nonlinear regressions tasks that is monotonically smooth, with a non-null gradient at the negative region, is developed. The proposed ac.f. is termed Leaky Exponential Linear Unit (or Leaky ELU, or LELU). The non-null gradient is proposed as a means to control an ad hoc score of the flexibility of an ac.f. and is a trainable parameter. Available nonlinear ac.f.’s have generally high flexibility scores which seem to be associated with proneness to overfitting. A null flexibility score can only represent linear data, on the other hand.
Smoothness is sought to avoid coding discontinuities at the predictions of the trained model. This is shown to be a result of the ReLU and Leaky ReLU ac.f.’s. The Leaky ReLU however has a trainable flexibility and this is shown to support the claim of beneficial and robust behaviour of this property for nonlinear regression tasks.
The results with the LELU are promising for highly nonlinear regressions tasks and large models. Its properties implicitly produce a behaviour similar to regularisation without strong penalisation of the training errors. The LELU does not show evidence of trading off training error with the overfitting/testing metric, at least not as strongly as other ac.f.’s. It also shows less sensitivity with model size and training points for nonlinear regression tasks.
@@
The author expects the proposed methods to be generalisable to other machine learning or dataset strategies. The observed properties of the proposed LELU should reduce the workload to select a good model setup and is as a such a highly desirable behaviour. The Laplace operator with its measure of the diffusion of a system provides a powerful tool for detection (and prevention) of overfitting in data regression. They should provide a sure hand in the process of implementing a model by bringing more robustness and clearer metrics.

### v2

@@ 4.1.3 Exponential Function and a Single Neuron
The previous problem is reduced to a set of three training points, $x_{i}=(-1,0,1)$, which a single neuron shall learn. This simple exercise can be performed by hand so two of the three points can be exactly matched. The obtained neuron parameters are used to infer a function as shown in Fig. 7, plotted along the ac.f. derivatives and training errors.
@@
From the perspective of learning nonlinear regression tasks, such weak signal is less then ideal. The gradient and Hessian of the loss can be computed by hand for this simple problem (motivated by [28]). This operation shows that the weak ac.fs. produce a vanishing loss gradient and higher conditioned Hessians, which is indicative of the tendency to poor learning and overfitting. The LELU on the other hand produces a clear gradient and a lower conditioned Hessian. With this simple exercise, the author intends to anecdotally show that the higher “flexibility” of the ac.f. is incurring in a weak signal for subsequent neurons and is a potential cause of instability showing itself as overfitting.
@@ 4.1.4 Other Findings with the Unidimensional problems
It is interesting to note that some ac.fs. struggle around $x=0$ in both case studies, even when this point is not a training data. A thorough explanation is challenging for a large model [24], but the author believes this can be partially understood by a similar analysis of the gradient as in the previous subsection. The loss-gradient is proportional to the summation of the gradient of the ac.f. times the training point coordinate. If the training point coordinate tends to zero, or adjacent training points cross $x=0$, this can further induce the loss gradient summation to vanish or “flicker”, and to destabilize the learning on an already unstable model setup. This behaviour is confirmed with a simple shift in $x$ by two units, as shown in Fig. 8, in which the local instability around $x=0$ is removed for the unstable models. Note that SiLU actually lost performance and presents overfitting, opposed to the behaviour shown in Fig. 5a, which corroborates the borderline stability of the model.
Another observation is that C-1 discontinuous ac.fs. generate “polyline” trained models. In these and some other setups not shown for conciseness, (leaky) ReLUs rather often default to piecewise linear connection between the training points or at least connect them with several segments of polyline. Such patterns are shown in the results in Figs. 5, 6 and 8, as well for the motor model to be discussed in what follows. Interestingly, in none of the cited ReLU-related literature in the introduction, the herein reported “polyline” pattern is commented.
The adopted canonical problems indicate that classical ac.fs. lack robustness with the problem setup for nonlinear regressions tasks when trained by large models. The proposition of the objectives in Sect. 2 for the LELU construction aims at mitigating the reported issues. LELU is able to achieve adequate results without heavily incurring in such penalties for the presented canonical problems. It can robustly learn the nonlinear regression task without memorising the training data, at the expense of slightly higher training MAEs, resembling an inherent but consistent auto-regularisation property of the ac.f. This behaviour is further confirmed on the more complex model discussed in what follows.
@@ 4.2 Electric Motor Model
The dataset is normalised so as to keep unitary mesh spacing in all dimensions to avoid cell skewness (which also favours the computation of the diffusion differentials), and kept in the positive range to support the more sensitive ac.fs. For the current study the trained data must exactly reproduce the noise-free training information [1]. A noiseless training is generally more challenging and is an additional reason for this choice in the current exercise.
A numerical experiment with the implemented [^3] Source files available in http://github.com/ebiga/gpr_aniso_trials NN with combinations of depth $n=[6,7,8]$ and width $m=[80,120,240,360]$ (fixed per layer) is executed. The model is trained for 9000 epochs with batch size 35. The learning rate is scheduled in the range [1e3,1e5][1e-3,1e-5] (5e45e-4 for the models with width $360$ and the $8\times 240$ case) with factor 0.5, patience 250, and cooldown 50. ELU, SiLU, and LELU ($\beta=0.4$) are run with 4 seeds of He-Normal and 1 seed of Glorot-Uniform weight initialisers for the $7\times 120$ and $8\times 120$ cases, and the standard deviations are presented as error bars. Where applicable, experiments with L2 regularisation are also reported.
@@
Figure 10: Electric motor model losses for different ac.fs. and model sizes indexed by $m^{n}/1e12$.
Figure 10 shows the MAE and diffusion metrics for the different model sizes and ac.fs. The tendency to lose generalisation shows as a lack of “convergence” of the losses with the increase in model size. For this complex problem the mean errors do not show large differences between the “good” and “bad,” so visual inspection guided by the results of the diffusion loss is necessary.
In overall, similar behaviours as with the canonical problems are also observed in this case. In special, the most promising classical ac.fs. SiLU and Softplus also present poor performance, consistent with the slice plots in Figs. 11 and 12. LELU successfully achieves lower losses amongst the tested ac.fs. and is also more robust to model size. As expected, the losses with different values of $\beta$ can be sensitive to model parameters. As an example, $\beta=0.2$ presents better results for smaller models and loses performance with larger ones, whereas $\beta=0.6$ presents the opposite trend. LELU with $\beta=0.4$ presents almost iso-performance with model parameters, and the sensitivity (error bars) to initialisation is also the smallest. In all cases, the results are more promising than any other ac.f.
@@
Figure 11 shows slices along $param2$, which is the fast-changing variable. The tanh and ReLU results are not shown due to poor performance. The region $param2=[0,2]$, where the highest slopes are seen, presents a tough task for the ac.fs. Considerable “localised” overfitting is observed for the ELU, SiLU and Softplus. Interestingly, the less flexible leaky ac.fs. show lower overfitting content and also less dependency on the model size. This nice behaviour prompted the current effort into assembling this property into a new ac.f. The Leaky ReLU results are “kinky” though and such behaviour is not ideal. The proposed LELU with $\beta=0.4$ shows a robust performance, indicating that the lower flexibility and smoothness are indeed advantageous.
@@
Figure 12 shows similar slices along the tight- and slow-changing $param3$. This slice allows for interesting behaviours to be more closely assessed, i.e., the spread of the ac.f. with different model sizes and the capacity to code small values. The less flexible leaky ac.fs. clearly show tighter and more representative capacity. SiLU and Softplus are also able to demonstrate a tighter spread, albeit with an offset error to the training data, whereas ELU is generally poorly behaved.
By inspection of ac.f. graphs in Fig. 1, it can be seen that both SiLU and Softplus present a wider range of slowly changing gradient, or in other words, ELU has a sharper gradient change. In the SiLU case, the gradient is actually never zero in the range plotted in Fig. 1. This observation has further prompted the effort into coding this property into the LELU. As seen in Fig. 1, this ac.f. slowly changes slopes and is therefore smoother than other ac.fs., thereby enabling a stronger a priori fitness for generalisable nonlinear regressions.
@@
Upon careful visual inspection of each NN result (see Fig. 13), depth 6 is not enough to capture all features, whereas depth 7 is able to do so but the MAE still indicates some distance (error) to the training points. Generally at least width 120 is necessary for acceptable results.
@@
Figure 14 shows slices with L2 (strength 1e41e-4) regularisation to reduce the overfitting tendency of ReLU and SiLU. As corroborated by the results in Fig. 10, the diffusion loss indicates a reduction in overfitting at the expense of an increase in the training MAE by more than a factor of 6. The performance does not improve over the LELU ($\beta=0.4$) and ReLU still keeps its “kinky” piecewise-linear predictions.
Another strategy is to implement a change of variables, for instance a power correction $\tilde{y}=y^{0.1}$, to reduce the large differences in the max and min values. A model trained on this function learns faster and easier, however the resulting errors when brought back to the original scale are also very large, which defeats the purpose. As already observed with the canonical problems, the LELU seems to provide adequate results without those drawbacks and extra steps.
@@ 5 Concluding Remarks
A parametric activation function for nonlinear regressions tasks that is monotonically smooth, with a non-null gradient at the negative region, is developed. The proposed ac.f. is termed Leaky Exponential Linear Unit (or Leaky ELU, or LELU). The non-null gradient is proposed as a means to control an ad hoc score of the flexibility of an ac.f. and is a trainable parameter. Available nonlinear ac.fs. have generally high flexibility scores which seem to be associated with proneness to overfitting. A null flexibility score can only represent linear data, on the other hand.
Smoothness is sought to avoid coding discontinuities at the predictions of the trained model. This is shown to be a result of the ReLU and Leaky ReLU ac.fs. The Leaky ReLU however has a trainable flexibility and this is shown to support the claim of beneficial and robust behaviour of this property for nonlinear regression tasks.
The results with the LELU are promising for highly nonlinear regressions tasks and large models. Its properties implicitly produce a behaviour similar to regularisation without strong penalisation of the training errors. The LELU does not show evidence of trading off training error with the overfitting/testing metric, at least not as strongly as other ac.fs. It also shows less sensitivity with model size and training sample size for nonlinear regression tasks. The parameter $\beta$ is proposed so it can be dedicatedly trained for better generalisation for different problems.
@@
The author expects the proposed methods to be generalisable to other machine learning or dataset strategies. The observed properties of the proposed LELU should reduce the workload to select a good model setup and is as such a highly desirable behaviour. The Laplace operator with its measure of the diffusion of a system provides a powerful tool for detection (and prevention) of overfitting in data regression. They should provide a sure hand in the process of implementing a model by bringing more robustness and clearer metrics.